OpenStudy (anonymous):

If the equation for a line is y > -x + 1 would I go down one back one on the graph? Or down one over one to the right?

jimthompson5910 (jim_thompson5910):

the slope is -1 so -1 = -1/1 = rise/run which means rise = -1 and run = 1 the rise of -1 means we go down 1 the run of +1 means you go to the right 1

OpenStudy (shadowlegendx):

Good job @jim_thompson5910